jabez177 (jabez177):

Who brought tobacco back from the New World? A. Francis Drake B. Sir Walter Raleigh C. Edmund Spenser D. William Shakespeare

OpenStudy (confluxepic):

Sir Walter Raleigh brought tobacco back from the New World.

OpenStudy (confluxepic):

The answer is B.

jabez177 (jabez177):

it is A i think

OpenStudy (confluxepic):

Sir Walter Raleigh is credited with taking the first "Virginia" tobacco to Europe, referring to it as tobah as early as 1578. In 1595 Anthony Chute published Tabaco, which repeated earlier arguments about the benefits of the plant and emphasised the health-giving properties of pipe-smoking.

OpenStudy (anonymous):

\(\Huge\mathbb{B}\)

OpenStudy (confluxepic):

That is what I said.

OpenStudy (anonymous):

I said it first. medal me

OpenStudy (confluxepic):

I said it first.

OpenStudy (confluxepic):

@jabez177

OpenStudy (anonymous):

\(\Huge\mathbb{COOLERTHANYOU}\)

OpenStudy (confluxepic):

\[\Huge\mathbb{YOU~WISH}\]

OpenStudy (anonymous):

@jabez177 Thats a really nice lambo, have you seen the 2016 sketches they look sick.